OpenStudy (anonymous):

Which of the following is the slope between the points (–2, 4) and (2, –2)(4 points) -3/2 -2/3 3/2 2/3 Thanks!

OpenStudy (anonymous):

Use this formula : \[Slope = \frac{y_2 -y_1}{x_2 -x_1}\]

OpenStudy (anonymous):

-6/4=-3/2

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

Thanks fazee!

OpenStudy (anonymous):

does anyone mind helping me with 2 more?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but give medals too plz

OpenStudy (anonymous):

So, the first is:

OpenStudy (anonymous):

what the unit and lesson ill see if ive done it already and give you them

OpenStudy (anonymous):

the slope between the points (0, –7) and (4, 12) is...? 19/4 -19/4 4/19 -4/19

OpenStudy (anonymous):

It's all about slope, that's my biggest problem

OpenStudy (anonymous):

And the other is Identify the correct x intercept and y intercept of the equation x + 7y = –21

OpenStudy (anonymous):

19/4

OpenStudy (anonymous):

Thanks! and the other is?

OpenStudy (anonymous):

for x int put y=0 x=-21

OpenStudy (anonymous):

and for y put x=0 y=-21/7

OpenStudy (anonymous):

Thanks alot!

OpenStudy (anonymous):

medal medal

OpenStudy (anonymous):

plz plz

OpenStudy (anonymous):

Okay, how do I do that?

OpenStudy (anonymous):

i want to show them to my girl friend that's simple just click on best response blue one

OpenStudy (anonymous):

There

OpenStudy (anonymous):

a and b